Circular No. 6512

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bw IN NGC 664 W. Li, Q. Qiao, Y. Qiu, and J. Hu, Beijing Astronomical Observatory

(BAO), report the discovery by the BAO Supernova Survey of a supernova located 19′′.7 west and 3′′.3 south of the center of NGC 664 (α = 1^{h}43^{m}46^{s}, δ = +4^{o}13′.3, equinox 2000.0). The following magnitudes of SN 1996bw were derived from unfiltered CCD images taken with the 0.60-m BAO reflector: Nov. 30, ∼ 17.5; Dec. 2, 17.2. Unfiltered CCD images of the same field (limiting mag ∼ 19.0) taken on Nov. 15 and 26 show no star at the position of SN 1996bw. A low-resolution, low-signal-to-noise spectrogram (0.96 nm/pixel, range 350–900 nm), taken with the BAO 2.16-m telescope by Li, J. Wei, and D. Xiu on Dec. 2.50 UT, suggests that this may be a type-Ib supernova prior to optical maximum; the spectrum is very blue, with no clear classification features. The overall spectrum is similar to that of SN 1984L (a prototype type-Ib supernova) at earlier times, but far bluer. There is conspicuous He I 570-nm absorption at 579 nm and Fe II absorption at 480 nm. The expansion velocity of the SN envelope is ∼ 10 500 km/s, derived from the He I 570-nm feature. The redshift of the galaxy (derived from the narrow Hα emission of the H II region) is 5563 km/s.

SGR 1806–20 K. Hurley, Space Sciences Laboratory, University of California, Berke-

ley, on behalf of the Ulysses Gamma-Ray Burst Team; C. Kouveliotou, Universities Space Research Association and Marshall Space Flight Center (MSFC), NASA; G. J. Fishman and C. A. Meegan, MSFC; and J. van Paradijs, University of Amsterdam and University of Alabama, Huntsville, report: “We have derived a much-refined position for the transient softγ-ray repeater reported on IAUC 6501, via triangulation between Ulysses and BATSE of two bursts on Nov. 19. The position is an annulus whose radius is 76^{o}.990, whose halfwidth is 0^{o}.0092 (3σ confidence level), centered at α = 23^{h}55^{m}34^{s}.9, δ = −30^{o}55′55′′ (equinox 2000.0). The combined triangulation annuli and BATSE error circles give error boxes whose areas are more than a factor of 100 smaller than the BATSE error circles alone. They are consistent with the known position of SGR 1806–20 (Murakami et al. 1994, Nature 368, 127), and the two annuli midlines pass within 4′′ of this position. We conclude that the source of these bursts, as well as of the ones that have occured since Oct. 30, is indeed SGR 1806–20. The latest short, soft event from this source was detected with BATSE on Nov. 23, so continued monitoring at other wavelengths is strongly encouraged.”

1996 December 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green